y=(x-1)√x²-2x+2

1
Expert's answer
2022-03-19T02:40:23-0400

Solution

"y=(x-1)\sqrt{x^2-2x+2}"


"\frac{dy}{dx}=(x-1)\frac{d}{dx}\sqrt{x^2-2x+2}+sqrt{x^2-2x+2}\frac{d}{dx}(x-1)"


"\frac{dy}{dx}=(x-1)\frac{1}{2}(x^2-2x+2)^{-\frac{1}{2}}\frac{d}{dx}(x^2-2x+2)+\sqrt{x^2-2x+2}(1-0)"


"\frac{dy}{dx}=(x-1)\frac{1}{2\sqrt{x^2-2x+2}}(2x-2)+\sqrt{x^2-2x+2}"


"\frac{dy}{dx}=(x-1)\frac{1}{2\sqrt{x^2-2x+2}}2(x-1)+\sqrt{x^2-2x+2}"


"\frac{dy}{dx}=\frac{(x-1)^2}{\sqrt{x^2-2x+2}}+\sqrt{x^2-2x+2}"


"\frac{dy}{dx}=\frac{(x-1)^2+\sqrt{x^2-2x+2}\sqrt{x^2-2x+2}}{\sqrt{x^2-2x+2}}"


"\frac{dy}{dx}=\frac{(x^2-2x+1)+(x^2-2x+2)}{\sqrt{x^2-2x+2}}"


"\frac{dy}{dx}=\frac{(2x^2-4x+3)}{\sqrt{x^2-2x+2}}"